A body is thrown vertically upward with a velocity u. The distance travelled by it in the 7th and 8th seconds are equal. The displacement in 8th seconds is equal yo ( take g= 10 m/s²

Grade:11

2 Answers

Vikas TU
7 years ago
Siply write the Newton’s second law of eqn. for each second.
For 7th second,
s7 = 7*u – 0.5*g*49.................(1)
s8 = 8*u – 0.5*g*64.............(2)
s6 = u*6 – 0.5*g*36..............(3)
 For 8th second,
Subtract the eqns. 2nd and 1st.
Hence.
Anurag
5 years ago
Time to reach maximum height is 7 second.
Displacement in the first second of its descent will be 1 /2 *g*1 = 5m. Distance in first second of descent is 5 m.